# From Demonstrations to Safe Deployment: Path-Consistent Safety Filtering for Diffusion Policies

### v1

@@ I Introduction
One of the main goals of robotics is to reduce the mental and physical burden on human workers by automating tedious and repetitive tasks. Recently, imitation learning using diffusion models [ho2020denoising] and flow matching [lipman2022flow] has enabled robots to successfully perform these complex, long-horizon manipulation tasks [chi2023diffusion, reuss2024multimodal]. These advances have been reinforced by the availability of large-scale demonstration datasets [o2024open] and the emergence of vision-language-action models [black2024pi_0, shukor2025smolvla]. These policies, which we collectively refer to as diffusion policies (DPs) in this work, can perform tasks in dynamic environments [ye2025ra, wang2025one]. However, since DPs are black-box models, they lack the required safety guarantees for avoiding collisions or safely interacting with dynamic objects. This prevents their deployment in human-centric environments, where safety has to be formally guaranteed.
Figure 1: Deploying DPs in dynamic environments with moving objects requires safeguarding mechanisms, as the intended policy actions may be unsafe. Reactive strategies, such as control barrier functions, often drive the agent into out-of-distribution (OOD) states not seen during training, leading to unpredictable behavior. We propose that safety mechanisms for DPs should remain consistent with the robot’s intended path to avoid out-of-distribution states and preserve high task success rates.
Many works have shown how safety can be ensured for learning-based policies during deployment [brunke2022safe, krasowski2023provably], for example, using control barrier functions [ames_2019_ControlBarrier, singletary2022safety] or predictive safety filters [wabersich_2021_PredictiveSafety, tearle2021predictive]. Most of these methods are reactive safety mechanisms that satisfy safety constraints by changing the path of the robot away from the dynamic object. However, the vast amount of available demonstrations [o2024open] is not recorded with any of such safety mechanisms in place. Therefore, as illustrated in Figure 1, reactive safety mechanisms often drive the system away from the demonstration data distribution into unseen states. DPs are particularly prone to errors in these out-of-distribution cases, increasing the risk of task failure [agia2025unpacking, xu2025can, romer2025failure]. To avoid this out-of-distribution situation, the interventions of a safety mechanism for DPs should ideally keep the system close to the desired actions of the policy.
We achieve this by proposing a Path-consistent safety filter (PACS) for DPs. The core idea is to use path-consistent braking to guarantee safety while keeping the system on its intended path. For this purpose, PACS translates the entire action sequence (chunk) generated by the policy into an intended trajectory, along which the robot is slowed down or stopped. We use set-based reachability analysis [althoff_2019_Effortless] to verify the safety constraints, allowing for real-time deployment at $1\text{\,}\mathrm{kHz}$.
@@ II Related Work
Deploying autonomous agents in safety-critical applications requires mechanisms to prevent harmful behavior. A common approach is to apply safety filters, which have been extensively reviewed in recent surveys [brunke2022safe, krasowski2023provably, wabersich_2023_DataDrivenSafety, hsu_2024_SafetyFilter]. Safety filters typically operate by verifying whether a failsafe action exists that can bring the robot to a safe state after executing the action proposed by the policy [hsu_2024_SafetyFilter]. Several techniques follow this approach: Model predictive safety filters [gros_2020_SafeReinforcement, wabersich_2021_PredictiveSafety] and control barrier functions [ames_2019_ControlBarrier, marvi_2021_SafeReinforcement, singletary2022safety, morton_2025_SafeTaskconsistent] formulate an optimization problem to compute the closest safe control input to the desired input. In contrast, methods based on reachability analysis [althoff_2015_IntroductionCORA, akametalu_2014_ReachabilitybasedSafe, fisac_2019_GeneralSafety, krasowski_2020_SafeReinforcement, shao_2021_ReachabilitybasedTrajectory, thumm_2022_ProvablySafe, thumm_2025_GeneralSafety] check whether the system remains outside unsafe states during the execution of the action and failsafe trajectory. These failsafe trajectories can be path-consistent, stopping the robot along its intended path [shao_2021_ReachabilitybasedTrajectory, thumm_2022_ProvablySafe], or non-path-consistent, guiding the robot away from unsafe regions.
@@
Compared to more classical programming-by-demonstration approaches [zeestraten2016online], the stochastic and end-to-end nature of diffusion policies makes them particularly prone to generating unsafe actions and behaving erratically, especially in unseen situations [agia2025unpacking, xu2025can, romer2025failure]. One way to address this problem is by adding an additional safety layer to the learned policy, for example, using control barrier functions [brunke2025semantically]. Another example is RAIL [jung2024railreachabilityaidedimitationlearning], which maintains a failsafe plan based on reachability analysis, defined as a stopping motion that stays as close as possible to the nominal trajectory. While theoretically well-grounded, such post-processing methods only consider static constraints so far and significantly distort the action distribution, leading to a notable drop in task performance [jung2024railreachabilityaidedimitationlearning]. To address this problem, several works directly incorporate safety constraints into the iterative action generation (denoising) process of diffusion- and flow-based policies. Adding cost gradients [carvalho2025motion, saha2024edmp] or classifier-free guidance [nikken2024denoising] can steer the generated action sequence towards safe regions, but these methods do not guarantee hard constraint satisfaction. As an alternative, iterative projections to a safe set [romer2024safe, romer2025diffusion, liang2025simultaneous] or control barrier functions [xiao2023safediffuser, dai2025safeflowmatchingrobot] have been injected into the denoising process to provide certain safety guarantees. However, these approaches still change the policy, and their high computational costs limit their applicability to low-dimensional systems [romer2024safe, romer2025diffusion] or offline planning [liang2025simultaneous, xiao2023safediffuser, dai2025safeflowmatchingrobot]. In summary, real-time safeguarding of DPs in dynamic environments without compromising task success remains an open problem.
@@ III Problem Statement
We aim to ensure a safe deployment of DPs when operating in close proximity to or interacting with dynamic objects, including humans. More specifically, we consider a robotic system with state $\bm{x}(t)\in\mathcal{X}$ and input $\bm{u}(t)\in\mathcal{U}$ at time $t$ that should perform a given task while adhering to a set of safety constraints. We assume that the state $\bm{x}$ of the system contains the pose of dynamic objects up to a bounded measurement error and that the objects have a bounded input set, e.g., a maximal velocity. We assume the availability of a set of trajectories demonstrated by experts $\mathcal{D}=\{\bm{o}_{k},\bm{a}_{k}\}_{k=0}^{K}$} in the form of observation-action pairs. The observations $\bm{o}\in\mathcal{O}$ can contain camera images, language commands, robot joint angles, and other environment information. The action $\bm{a}\in\mathcal{A}$ is a high-level goal that the robot should reach, e.g., a desired joint configuration. From the expert demonstrations, a DP $\pi:\mathcal{A}^{H}\times\mathcal{O}\rightarrow\mathbb{R}_{+}$ is trained that returns a distribution over action chunks $\bm{A}\in\mathcal{A}^{H}$ of length $H$. In this work, we aim to design a controller $\bm{k}:\mathcal{X}\times\mathcal{A}^{H}\times\mathbb{R}\rightarrow\mathcal{U}$} that takes as input the latest action chunk $\bm{A}$ generated by the policy and returns a safe control input $\bm{u}_{\text{safe}}(t)$.
@@ IV Preliminaries
This section describes the technical background on DPs and reachability-based safety verification.
@@ IV-A Imitation Learning with Diffusion Policies
The demonstration data $\mathcal{D}$ encodes the desired behavior for solving the task. A DP $\pi$ is trained to approximate the conditional action distribution in the demonstrations [chi2023diffusion]. Once trained, diffusion- and flow-based policies generate new actions by iteratively transforming a random noise sample into the learned target distribution [ho2020denoising, lipman2022flow]. Most DPs return a chunk of $H$ consecutive actions $\bm{A}_{k}=(\bm{a}_{k},\bm{a}_{k+1},\dots,\bm{a}_{k+H-1})$} instead of only the next action to improve the temporal consistency and execution speed [chi2023diffusion, zhao2023learning]. During deployment, we can iteratively sample new action chunks $\bm{A}_{k}\sim\pi(\cdot\,|\,\bm{o}_{k})$}, apply the first $h\leq H$ actions $\bm{a}_{k:k+h-1}$ to the system for a time $\Delta t$, respectively, and replan at timestep k+h{k+h}, i.e., after a time $h\Delta t$ has passed. We refer to $H$ as the action chunk length and $h$ as the action execution steps. This general formulation of imitation learning policies encompasses state-of-the-art generative single-task policies [chi2023diffusion, reuss2024multimodal] and generalist vision-language-action models [black2024pi_0, shukor2025smolvla].
@@ IV-B Safety Constraints for Dynamic Objects
The robot and the objects are considered as dynamic systems, and we define the trajectory of a system at time $t$ for a given initial state $\bm{x}_{0}\in\mathcal{X}_{0}$} and a possible input trajectory $\bm{u}(\cdot)$} by $\chi(t;\bm{x}_{0};\bm{u}(\cdot))$}. Following [althoff_2019_Effortless, p. 5], we formally introduce the reachable set $\mathcal{R}(t)$} of a system at time $t$ for a compact set of possible inputs $\mathcal{U}$ by
@@
and denote all states that are reachable during a time interval as $\mathcal{R}([t_{a},t_{b}])=\bigcup_{t\in[t_{a},t_{b}]}\mathcal{R}(t)$}. We denote the reachable occupancy by $\mathcal{C}(t)$. For the reachable occupancy computation of the robot $\mathcal{C}_{\text{robot}}(t)$, we assume a bounded tracking error of the desired trajectory. This assumption could be replaced by applying more advanced methods as presented in [stefan2025ControllerSynthesis].
Following prior work [althoff_2019_Effortless, liu_2021_OnlineVerification, beckert_2017_multiple_safety], we define a collision between an object and the robot if their reachable occupancies $\mathcal{C}_{\text{object}}(t)$ and $\mathcal{C}_{\text{robot}}(t)$ intersect, i.e., a collision event is defined as
@@
Depending on the task and object, collisions should either be avoided altogether or restricted to non-harmful, low-force interactions. Accordingly, we define two types of safety constraints in accordance with ISO/TS 15066:2016. Power and force limiting (PFL) targets human-robot collaboration with close interaction and allows contacts with humans as long as the kinetic energy of the robot is below pain and injury thresholds $T_{\text{safe}}(t)$. So, the PFL safety constraint is
@@
The safe energy threshold $T_{\text{safe}}(t)$ can depend on the collision type (constrained collisions for clamping or unconstrained collisions for free movement), the shape of the robot part in contact (e.g., sharp or blunt edges), or the human body part affected in the worst case [kirschner_2024_constrained, kirschner_2024_unconstrained, thumm_2025_GeneralSafety]. In contrast, speed and separation monitoring (SSM) targets coexistence without close interaction, as it does not permit any contact with humans. If a collision can occur, the robot must be at a complete stop with zero kinetic energy, i.e., $T_{\text{safe}}(t)=0$ in (3).
V Path-Consistent Safety Filter for Diffusion Policies
@@ V-A Safety-Permitting Diffusion Policy
To maintain high task performance despite path-consistent safety interventions, we need to design the policy in such a way that slowing down or stopping the robot does not in itself represent an out-of-distribution situation. To achieve this, we only include RGB images $\bm{I}(t)$ and joint angles $\bm{q}(t)$ into the observation $\bm{o}(t)$} and no explicit velocity information, which is a common approach in recent works [black2024pi_0, shukor2025smolvla]. Our approach is applicable to various action spaces, but we restrict our presentation to joint velocity control for clarity. In this case, the actions can be the delta joint positions $\Delta\bm{\hat{q}}_{i}=(\bm{\hat{q}}_{i+1}-\bm{\hat{q}}_{i})$, and an action chunk is given by (Algorithm 1, line 5)
$\bm{A}(t)=(\Delta\bm{\hat{q}}_{0},\dots,\Delta\bm{\hat{q}}_{H-1})\sim\pi(\cdot\,|\,\bm{o}(t)).$ (4)
Here, we use $\hat{\cdot}$ to distinguish the desired actions generated by the policy from the actual safe commands sent to the robot. By integrating the desired joint positions forward in time, starting from the current state $\bm{\hat{q}}_{0}$, we compute a sequence of waypoints
$\bm{\hat{p}}(t)=(\bm{\hat{q}}_{0},\dots,\bm{\hat{q}}_{h}),\quad\text{where}\quad\bm{\hat{q}}_{0}=\bm{q}(t),$ (5)
describing the desired path of the policy (Algorithm 1, line 6).
Algorithm 1 Path-Consistent Safety Filter (PACS) for diffusion policies (DPs) and vision-language-action models (VLAs).
@@
4:while $\neg\textit{success}$ do $\triangleright$ High-level (diffusion) steps
@@
7:  $(\bm{\hat{q}}_{0}=\bm{q}(t),\dots,\bm{\hat{q}}_{h})\leftarrow\textsc{Integrate}(\bm{A}(t))$
8:  $\chi_{I}([t,t_{h}])\leftarrow\textsc{ComputeTrajectory}((\bm{\hat{q}}_{0},\ldots,\bm{\hat{q}}_{h}))$
9:  $\textit{shield}\leftarrow\textsc{SetTrajectory}(\chi_{I}([t,t_{h}]))$
10:  for $k\leftarrow 0$ to $\frac{h\Delta t}{\alpha_{s}}$ do $\triangleright$ Low-level safety steps
@@ V-B Path-Consistent Safety for Diffusion Policies
To ensure that our safety filter does not steer the policy into out-of-distribution states, we adopt a path-consistent safety formulation. The key idea is to guarantee safety by modifying the joint speed $\dot{\bm{q}}(t)$ and acceleration $\ddot{\bm{q}}(t)$ while staying on the intended trajectory. Existing path-consistent safety filters for learning-based policies using reachability analysis [thumm_2025_GeneralSafety, thumm_2022_ProvablySafe] can only take as action input a single goal position with zero velocity. However, DPs and vision-language-action models predict a chunk (4) of multiple consecutive actions together and operate at a relatively low update frequency $1/(h\Delta t)$. Therefore, directly applying single-action approaches [thumm_2025_GeneralSafety, thumm_2022_ProvablySafe] to these policies either leads to a very slow execution by stopping at each waypoint or leaves out intermediate action steps, sacrificing fine-grained control and causing a deviation from the desired behavior.
@@
To construct smooth trajectories that align with the predicted action chunk, we propose an intended trajectory planning module (cf. Figure 2). For each action chunk (4), we compute a time-optimized intended trajectory $\chi_{I}([t,t_{h}])=\big\{\bm{x}(\tau)\,\big|\,\tau\in[t,t_{h}]\}$} (Algorithm 1, line 7), where $\bm{x}(t)=\big(\bm{q}(t),\dot{\bm{q}}(t),\ddot{\bm{q}}(t),\dddot{\bm{q}}(t)\big)$ is the state of the system, by solving the optimization problem
$\min_{t_{h},\,\chi_{I}([t,t_{h}])}\quad$ $t_{h}-t$ (6a)
s.t. $\forall\tau\in[t,t_{h}]:\;\bm{x}(\tau)\in\mathcal{X},$ (6b)
$\forall k\in\{1,\dots,h-1\}:$
$\qquad\exists\,t_{k}\in(t_{k-1},t_{k+1}):~\bm{q}(t_{k})=\bm{\hat{q}}_{k},$ (6c)
$\bm{q}(t_{0}=t)=\bm{\hat{q}}_{0},\qquad\bm{q}(t_{h})=\bm{\hat{q}}_{h},$ (6d)
@@
where the set $\mathcal{X}$ defines robot-specific kinematic and dynamic constraints $\bm{q}_{\min}\leq\bm{q}(t)\leq\bm{q}_{\max}$, $\|\dot{\bm{q}}(t)\|\leq\dot{q}_{\max}$, $\|\ddot{\bm{q}}(t)\|\leq\ddot{q}_{\max}$, and $\|\dddot{\bm{q}}(t)\|\leq\dddot{q}_{\max}$. Adjusting the velocity, acceleration, and jerk limits allows us to adapt the execution speed of the action chunk, which gives us a more fine-grained control over the execution speed of the policy compared to modifying the actions directly. Importantly, our intermediate planning module provides a trajectory (6) describing the desired behavior of the policy, which we forward to our path-consistent safety filter (Algorithm 1, line 8).
@@
The safety filter (Algorithm 1, line 9-14) operates at a high frequency and formally guarantees constraint satisfaction while staying on the desired path (5). Based on the current state of the robot $\bm{x}_{0}=\bm{x}(t)$ and the measured object positions $\bm{m}_{\text{o}}(t)$ (Algorithm 1, line 10), we compute a path-consistent monitored trajectory [thumm_2025_GeneralSafety] with intended step $t_{0},t_{\text{I}}]$ and failsafe trajectory (stopping motion) $t_{\text{I}},t_{\text{F}}]$
$\chi_{\text{M}}(t)=\begin{cases}\chi_{I}\big(t;\,\bm{x}_{0},\bm{u}_{I}(\cdot)\big),&t\in[t_{0},t_{\text{I}}],\\[3.99994pt] \chi_{F}\big(t;\,\bm{x}_{I},\bm{u}_{F}(\cdot)\big),&t\in[t_{\text{I}},t_{\text{F}}].\end{cases}$ (7)
@@
At each low-level safety timestep of duration $\alpha_{\text{s}}$, we verify the monitored trajectory (7) for the task-specific safety constraints as described in Section IV-B. If the verification is successful, we execute the intended trajectory $\bm{u}_{\text{safe}}(t)=\bm{u}_{\text{I}}(t)$}; otherwise, we proceed one step on the last successfully verified failsafe trajectory, i.e., $\bm{u}_{\text{safe}}(t)=\bm{u}_{\text{F}}(t)$}. This process is repeated until we receive a new action chunk (4) from the policy after $h\Delta t/\alpha_{\text{s}}$ safety steps. Since this safeguarding mechanism is provably safe by induction, as shown in prior work [thumm_2025_GeneralSafety], PACS enables the safe deployment of DPs.
@@
(b) Handover: Grab Lego blocks from a hand and put them into the box.
@@
Figure 3: Visualization of our three real-world tasks, which require safe and reactive motion in close proximity to the human body. Sorting represents a coexistence task where no collisions are permitted, whereas Handover and Feeding are collaboration tasks requiring non-harmful, low-force interactions.

### v2

@@ I Introduction
One of the main goals of robotics is to reduce the mental and physical burden on human workers by automating tedious and repetitive tasks. Recently, imitation learning using diffusion models [16] and flow matching [25] has enabled robots to successfully perform these complex, long-horizon manipulation tasks [12, 34]. These advances have been reinforced by the availability of large-scale demonstration datasets [32] and the emergence of vision-language-action models [8, 40]. Diffusion policies (DPs), i.e., robot policies parameterized by generative diffusion or flow-matching models, can perform tasks in dynamic environments [51, 48]. However, since DPs are black-box models, they lack safety guarantees for avoiding collisions or safely interacting with dynamic objects. This prevents their deployment in human-centric environments, where safety has to be formally guaranteed.
Figure 1: Deploying DPs in dynamic environments with moving objects requires safeguarding mechanisms, as the intended policy actions may be unsafe. Reactive strategies, such as control barrier functions, often drive the agent into out-of-distribution (OOD) states not seen during training, leading to unpredictable behavior. We propose that safety mechanisms for DPs should remain consistent with the intended path of the robot to avoid out-of-distribution states and preserve high task success rates.
Many works have shown how safety can be ensured for learning-based policies during deployment [9, 22], e.g., using control barrier functions [5, 41] or predictive safety filters [47, 42]. Most of these methods are reactive safety mechanisms that satisfy safety constraints by changing the path of the robot away from the dynamic object. However, the vast amount of available demonstrations [32] is not recorded with any of such safety mechanisms in place. Therefore, as illustrated in Figure 1, reactive safety mechanisms often drive the system away from the demonstration data distribution into unseen states. DPs are particularly prone to errors in these out-of-distribution cases, increasing the risk of task failure [1, 50, 36]. To avoid this out-of-distribution situation, the interventions of a safety mechanism for DPs should ideally keep the system close to the desired actions of the policy.
We achieve this by proposing a Path-consistent safety filter (PACS) for DPs. The core idea is to use path-consistent braking to guarantee safety while keeping the system on its intended path. To this end, PACS translates the entire action sequence (chunk) generated by the policy into an intended trajectory, along which the robot is slowed down or stopped. We use set-based reachability analysis [3] to verify safety constraints, enabling real-time deployment at $1\text{\,}\mathrm{kHz}$.
@@ II Related Work
Deploying autonomous agents in safety-critical applications requires mechanisms to prevent harmful behavior. A common approach is to apply safety filters, which have been extensively reviewed in recent surveys [9, 22, 46, 17]. Safety filters typically verify whether a failsafe action exists that can bring the robot to a safe state after executing the action proposed by the policy [17]. Several techniques follow this approach: Model predictive safety filters [15, 47] and control barrier functions [5, 29, 41, 30] formulate an optimization problem to compute the safe control input closest to the desired input. In contrast, methods based on reachability analysis [4, 2, 14, 23, 39, 43, 44] ensure safety by keeping available a provably correct failsafe trajectory. These failsafe trajectories can be path-consistent, stopping the robot along its intended path [52, 39, 43], or non-path-consistent, guiding the robot away from unsafe regions.
@@
Compared to more classical programming-by-demonstration approaches [53], the stochastic and end-to-end nature of diffusion policies makes them particularly prone to unsafe and erratic behavior, especially in unseen situations [1, 50, 36]. One way to address this problem is by adding an additional safety layer to the learned policy, e.g., using control barrier functions [10]. Another example is RAIL [19], which maintains a failsafe plan based on reachability analysis, defined as a stopping motion that stays as close as possible to the nominal trajectory. While theoretically well-grounded, such post-processing methods only consider static constraints so far and significantly distort the action distribution, leading to a notable drop in task performance [19]. To address this problem, several works directly incorporate safety constraints into the iterative action generation (denoising) process of diffusion- and flow-based policies. Adding cost gradients [11] or classifier-free guidance [31] can steer the generated action sequence towards safe regions, but these methods do not guarantee hard constraint satisfaction. As an alternative, iterative projections to a safe set [35, 37, 24] or control barrier functions [49, 18] have been injected into the denoising process to provide certain safety guarantees. However, these approaches still change the policy, and their high computational costs limit their applicability to low-dimensional systems [35, 37] or offline planning [24, 49, 18]. In summary, real-time safeguarding of DPs in dynamic environments without compromising task success remains an open problem.
@@ III Problem Statement
We aim to ensure a safe deployment of DPs when operating in close proximity to or interacting with dynamic objects, including humans. More specifically, we consider a robotic system with state $\bm{x}(t)\in\mathcal{X}$ and input $\bm{u}(t)\in\mathcal{U}$ at time $t$ that should perform a given task while adhering to a set of safety constraints. We assume that the state $\bm{x}$ of the system contains the pose of dynamic objects up to a bounded measurement error and that the objects have a bounded input set, e.g., a maximal velocity. We assume the availability of a set of trajectories demonstrated by experts $\mathcal{D}=\{\bm{o}_{k},\bm{a}_{k}\}_{k=0}^{K}$} in the form of observation-action pairs. The observations $\bm{o}\in\mathcal{O}$ may include robot joint angles, camera images, language commands, and other environmental information. The action $\bm{a}\in\mathcal{A}$} is a high-level goal that the robot should reach, e.g., a desired joint configuration. From the expert demonstrations, a DP $\pi:\mathcal{A}^{H}\times\mathcal{O}\rightarrow\mathbb{R}_{+}$ is trained that returns a distribution over action chunks $\bm{A}\in\mathcal{A}^{H}$ of length $H$. In this work, we aim to design a controller $\bm{k}:\mathcal{X}\times\mathcal{A}^{H}\times\mathbb{R}\rightarrow\mathcal{U}$} that takes as input the latest action chunk $\bm{A}$ generated by the policy and returns a safe control input $\bm{u}_{\text{safe}}(t)$.
@@ IV Preliminaries
This section describes the technical background on DPs and reachability-based safety verification. Throughout the paper, we use $(\cdot)$ as a placeholder for the arguments of a probability distribution or function.
@@ IV-A Imitation Learning with Diffusion Policies
The demonstration data $\mathcal{D}$ encodes the desired behavior for solving the task. A DP $\pi$ is trained to approximate the conditional action distribution in the demonstrations [12]. Once trained, DPs and VLAs generate new actions by iteratively transforming a random noise sample into the learned target distribution [16, 25]. Most DPs return a chunk of $H$ consecutive actions $\bm{A}_{k}=(\bm{a}_{k},\bm{a}_{k+1},\dots,\bm{a}_{k+H-1})$} instead of only the next action to improve the temporal consistency and execution speed [12, 54]. During deployment, we can iteratively sample new action chunks $\bm{A}_{k}\sim\pi(\cdot\,|\,\bm{o}_{k})$}, apply the first $h\leq H$ actions $\bm{a}_{k:k+h-1}$ to the system for a time $\Delta t$, respectively, and replan after a time $h\Delta t$ has passed. This general formulation of imitation learning policies encompasses state-of-the-art generative single-task policies [12, 34] and generalist vision-language-action models [8, 40].
@@ IV-B Safety Constraints for Dynamic Objects
The robot and the objects are considered as dynamic systems, and we define the trajectory of a system at time $t$ for a given initial state $\bm{x}_{0}\in\mathcal{X}_{0}$} and a possible input trajectory $\bm{u}(\cdot)$} by $\chi(t;\bm{x}_{0};\bm{u}(\cdot))$}. Following [3, p. 5], we formally introduce the reachable set $\mathcal{R}(t)$} of a system at time $t$ for a compact set of possible inputs $\mathcal{U}$ by
@@
and denote all states that are reachable during a time interval as $\mathcal{R}([t_{a},t_{b}])=\bigcup_{t\in[t_{a},t_{b}]}\mathcal{R}(t)$}. We denote the reachable occupancy by $\mathcal{C}(t)$. For the reachable occupancy computation of the robot $\mathcal{C}_{\text{robot}}(t)$, we assume a bounded tracking error of the desired trajectory. This assumption could be replaced by applying more advanced methods as presented in [27]. To compute the reachable occupancies of dynamic objects $\mathcal{C}_{\text{object}}(t)$, we assume access to measurements of positions of objects with bounded measurement errors. Moreover, we consider velocity and acceleration limits, and further include sensor delay and bounded measurement errors on position and velocity, following the approach in [3].
Following prior work [3, 26, 6], we define a collision between an object and the robot if their reachable occupancies $\mathcal{C}_{\text{object}}(t)$ and $\mathcal{C}_{\text{robot}}(t)$ intersect, i.e., a collision event is defined as
@@
Depending on the task and object, collisions should either be avoided altogether or restricted to non-harmful, low-force interactions. Accordingly, we define two types of safety constraints in accordance with ISO/TS 15066:2016. Power and force limiting (PFL) targets human-robot collaboration with close interaction and allows contacts with humans as long as the kinetic energy of the robot is below pain and injury thresholds $T_{\text{safe}}(t)$, resulting in the PFL safety constraint
@@
The safe energy threshold $T_{\text{safe}}(t)$ can depend on the collision type (constrained collisions for clamping or unconstrained collisions for free movement), the shape of the robot part in contact (e.g., sharp or blunt edges), or the human body part affected in the worst case [20, 21, 44]. In contrast, speed and separation monitoring (SSM) does not permit any contact with humans, targeting coexistence without close interaction. If a collision can occur, the robot must be at a complete stop with zero kinetic energy, i.e., $T_{\text{safe}}(t)=0$ in (3).
V Path-Consistent Safety Filter
@@ V-A Safety-Permitting Diffusion Policy
To maintain high task performance despite path-consistent safety interventions, we need to design the policy in such a way that slowing down or stopping the robot does not in itself represent an out-of-distribution situation. To achieve this, we only include RGB images $\bm{I}(t)$ and joint angles $\bm{q}(t)$ into the observation $\bm{o}(t)$} and no explicit velocity information, which is a common approach in recent works [8, 40]. Our approach is applicable to various action spaces, but we restrict our presentation to joint velocity control for clarity. In this work, the actions are the delta joint positions $\bm{a}_{i}=\Delta\bm{\hat{q}}_{i}=(\bm{\hat{q}}_{i+1}-\bm{\hat{q}}_{i})$} and we use $\hat{\cdot}$ to distinguish the desired actions generated by the policy from the actual safe commands sent to the robot. Without loss of generality, we reset the clock in each step of the planning algorithm and safety filter, i.e., $t_{0}=0$. By integrating the desired joint positions forward in time, starting from the current joint position $\bm{q}(0)$, we compute a sequence of waypoints
$\bm{\hat{p}}(0)=(\bm{\hat{q}}_{0},\dots,\bm{\hat{q}}_{h}),\quad\text{where}\quad\bm{\hat{q}}_{0}=\bm{q}(0),$ (4)
describing the desired path (Algorithm 1, line 6).
Algorithm 1 Path-consistent safety filter for diffusion policies (DPs) and vision-language-action models (VLAs).
@@
4:while $\neg\textit{success}$ do $\triangleright$ Policy steps
@@
7:  $(\bm{\hat{q}}_{0},\dots,\bm{\hat{q}}_{h})\leftarrow\textsc{Integrate}(\bm{A}(t))$
8:  $\chi_{I}([t_{0},t_{h}])\leftarrow\textsc{ComputeTrajectory}((\bm{\hat{q}}_{0},\ldots,\bm{\hat{q}}_{h}))$
9:  $\textit{shield}\leftarrow\textsc{SetTrajectory}(\chi_{I}([t_{0},t_{h}]))$
10:  for $k\leftarrow 0$ to $\frac{h\Delta t}{\alpha_{s}}$ do $\triangleright$ Safety steps
@@ V-B Path-Consistent Safety for Diffusion Policies
To ensure that our safety filter does not steer the policy into out-of-distribution states, we adopt a path-consistent safety formulation. The key idea is to guarantee safety by modifying the joint speed $\dot{\bm{q}}(t)$ and acceleration $\ddot{\bm{q}}(t)$ while staying on the intended trajectory. Existing path-consistent safety filters for learning-based policies using reachability analysis [44, 43] can only take as action input a single goal position with zero velocity. However, DPs and vision-language-action models predict a chunk $\bm{A}(t)$ of multiple consecutive actions together and operate at a relatively low update frequency $1/(h\Delta t)$. Therefore, directly applying single-action approaches [44, 43] to these policies either leads to a very slow execution by stopping at each waypoint or leaves out intermediate action steps, sacrificing fine-grained control and causing a deviation from the desired behavior.
@@
To construct smooth trajectories that align with the predicted action chunk, we propose an intended trajectory planning module (cf. Figure 2). For each action chunk $\bm{A}(t)$ starting from the initial time step $t_{0}=0$, we compute a time-optimized intended trajectory $\chi_{I}([0,t_{h}])=\big\{\bm{x}(\tau)\,\big|\,\tau\in[0,t_{h}]\}$} (Algorithm 1, line 7), where $\bm{x}(\tau)=\big(\bm{q}(\tau),\dot{\bm{q}}(\tau),\ddot{\bm{q}}(\tau),\dddot{\bm{q}}(\tau)\big)$ is the state of the system, by solving the optimization problem
$\min_{t_{h},\,\chi_{I}([0,t_{h}])}\quad$ $t_{h}$ (5a)
s.t. $\forall\tau\in[0,t_{h}]:\;\bm{x}(\tau)\in\mathcal{X},$ (5b)
$\forall k\in\{0,\dots,h\}:\bm{q}(t_{k})=\bm{\hat{q}}_{k},$ (5c)
@@
where the set $\mathcal{X}$ defines robot-specific kinematic and dynamic constraints $\bm{q}_{\min}\leq\bm{q}(\tau)\leq\bm{q}_{\max}$, $\|\dot{\bm{q}}(\tau)\|\leq\dot{q}_{\max}$, $\|\ddot{\bm{q}}(\tau)\|\leq\ddot{q}_{\max}$, and $\|\dddot{\bm{q}}(\tau)\|\leq\dddot{q}_{\max}$. Adjusting the velocity, acceleration, and jerk limits allows us to adapt the execution speed of the action chunk, which gives us a more fine-grained control over the execution speed of the policy compared to modifying the actions directly. Importantly, our intermediate planning module provides a trajectory by solving (5), which we forward to our path-consistent safety filter (Algorithm 1, line 8).
@@
The safety filter (Algorithm 1, line 9-14) operates at a high frequency and formally guarantees constraint satisfaction while staying on the desired path (4). Based on the current state of the robot $\bm{x}_{0}=\bm{x}(0)$ and the measured object positions $\bm{m}_{\text{o}}(t)$ (Algorithm 1, line 10), we compute a path-consistent monitored trajectory [44] by concatenating an intended trajectory (desired or recovery motion) $0,t_{\text{I}}]$ and a failsafe trajectory (stopping motion) $t_{\text{I}},t_{\text{F}}]$
$\chi_{\text{M}}(t)=\begin{cases}\chi_{I}\big(\tau;\,\bm{x}_{0},\bm{u}_{I}(\cdot)\big),&\tau\in[0,t_{\text{I}}],\\[3.99994pt] \chi_{F}\big(\tau;\,\bm{x}_{I},\bm{u}_{F}(\cdot)\big),&\tau\in[t_{\text{I}},t_{\text{F}}].\end{cases}$ (6)
@@
At each safety timestep of duration $\alpha_{\text{s}}$, we verify the monitored trajectory (6) for the task-specific safety constraints as described in Section IV-B. If the verification is successful, we execute the intended trajectory $\bm{u}_{\text{safe}}(t)=\bm{u}_{\text{I}}(t)$}; otherwise, we proceed with the last successfully verified failsafe trajectory that brings the system to an invariable safe set, i.e., $\bm{u}_{\text{safe}}(t)=\bm{u}_{\text{F}}(t)$}. We repeat this process until receiving a new action chunk from the policy after $h\Delta t/\alpha_{\text{s}}$ safety steps. Since this safeguarding mechanism is provably safe by induction, as shown in prior work [44], PACS enables the safe deployment of DPs.
@@
(b) Handover: Grab blocks from a moving hand and put them into the box.
@@
Figure 3: Visualization of our three real-world tasks, which require safe and reactive motion in close proximity to the human body. Sorting represents a coexistence task, where no collisions are permitted, whereas Handover and Feeding are collaboration tasks requiring non-harmful, low-force interactions.